OpenStudy (anonymous):

What is the solution set for |2x – 3| = 11?

OpenStudy (anonymous):

Firstly solve for 2x - 3 = 11 Add 3 both the sides and then divide it by 2 to get x..

OpenStudy (cwrw238):

the 2x - 3 in the | | can be positive or negative so 2x - 3 = 11 or 2x -3 = -11 solve these 2 equations to get the solution set

OpenStudy (anonymous):

-7 and 4?

OpenStudy (cwrw238):

not quite switch the signs around 2x - 3 = 11 dd 3 to both sides 2x = 14 x = 14/2 = 7

OpenStudy (anonymous):

The second equation formed will be: No no you are right my mistake..

OpenStudy (anonymous):

-4 and 7?

OpenStudy (anonymous):

You will get 7 and -4 @china314

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

You are right..